OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? please tell me the answer im running out of time

OpenStudy (igreen):

Welcome to Open Study! We can plug this into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{1+5}{4-1}\) Can you simplify that? @caitlyn9999

OpenStudy (igreen):

Can you add 1 + 5? @caitlyn9999

OpenStudy (bibby):

rofl the question below this is literally the exact same thing

OpenStudy (anonymous):

6/3

OpenStudy (igreen):

O_O

OpenStudy (igreen):

It is.. xD

OpenStudy (igreen):

Lol..nice picture @CivC97

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Not anymore.

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

r.i.p

OpenStudy (igreen):

@caitlyn9999 You there..? :l

OpenStudy (anonymous):

I ran out of time already , thanks anyways

OpenStudy (bibby):

try reviewing next time

OpenStudy (igreen):

^^^^^^^^^^^^^^^^^

OpenStudy (aaronandyson):

^_^